Note:

1 mole of a compound is equivalent to its molecular mass

1 mole of a compound is equal to 6.02x10^23 molecules

The percentage is a direct equivalent to grams

Answers should be in 3 significant figures

No need to include the units and the compound in the answers


  1. A 60.00 g sample of tetraethyllead, a gasoline additive, is found to contain 38.43 g lead, 17.83 g carbon, and 3.74 g hydrogen. Find its empirical formula. 
  2. A 170.00 g sample of an unidentified compound contains 29.84 g sodium, 67.49 g chromium, and 72.67 g oxygen. Find the empirical formula.  
  3. Determine the molecular formula of the compound with an empirical formula of CH and a formula mass of 78.110 u
1
Expert's answer
2021-06-30T03:56:38-0400

in a 60g sample of tetraethyl-lead, a gasoline is addictive , is found contain 38.43g lead , 17.83g carbon and 3.74g hydrogen.

percentage mass of Pb = 38.43/60 × 100

≈ 64.05 %

percentage mass of C = 17.83/60 × 100

≈ 29.71%

percentage mass of H = 3.74/60 × 100

≈ 6.23%

now, finding percentage - atomic mass ratio

Pb => 64.05/207 = 0.3094

C => 29.71/12 = 2.475

H => 6.23/1 = 6.23

finding simple ratio of above terms

Pb : C : H = 1 : 8 : 20

hence, empirical formula is "PbC_6H_{20}" or, "PbC_8H_{20}"






The amu of CH is the sum of the amu of an atom of C-12 and an atom of hydrogen which gives 12 + 1.00794 = 13.00794. The molecular mass of the compound we have to determine which is 78.110. Dividing this by 13.00794 gives us 6.0047.